# Bridging Today and the Future of Humanity: AI Safety in 2024 and Beyond

### v2

@@ 4.1 Rookie Level: Data-Driven Era
The Data-Driven Era represents the early phase of AI development, where AI models heavily rely on the quality, availability, and quantity of training data [131]. AI makes probabilistic inferences based on patterns found in massive training datasets, thus failing to achieve genuine understanding of data [96, 239]. While AI in this era cannot be regarded as real intelligence, it assists people to solve problems and enhances productivity in the society.
Stage 1: Past, Current, and the Near Future. This stage is characterized by diverse AI models for processing various forms of data across different scenarios [131, 50, 63, 210, 231, 260]. Researchers have been engaging in refining model structures and developing advanced algorithms, enabling more sophisticated analysis of data in different formats. The thrive of LLMs exemplifies this progress, showcasing unprecedented capabilities of AI models in language processing and generation [260, 286, 231]. Concurrently, we observe significant advancements in other fields such as computer vision [101, 221, 256], with progress in image processing [235, 25, 216] and video analysis [223, 261, 267], as well as advancements in audio recognition [64, 178, 4, 106]. This stage is driven by exploring vast datasets and the refining algorithms continuously, which consistently pushes the boundaries of AI capabilities within specific domains.
Stage 2: Interdisciplinary AI. This stage is characterized by leveraging AI to transform traditional fields, such as healthcare [231, 2], finance [260, 282], and education [184, 135], through cross-disciplinary applications in human society. Mature multi-agent systems [116, 246, 263, 79] that leverage AI entities with different specializations will be widely utilized to assist people and solve sophisticated problems. Such agents participate in decision-making processes, or represent users to finish some tasks, providing more sophisticated solutions of a given problem. In healthcare, for example, a multi-agent system may involve collaboration between various medical agent specialties and integrate data from diverse sources to formulate comprehensive treatment plans. The rise of interdisciplinary AI integration will amplify the utility of AI in more complex scenarios. Although AI at this stage is still not real intelligence, it becomes deeply integrated in different domains in human society, and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
@@ 4.2 Champion Level: Neural Augmentation Era
In this era, AI models have been widely applied in human society, and humans may witness breakthroughs in advanced technologies and fundamental theories, which may extend the boundary of human cognition and life experience.
Stage 3: Breakthroughs in Advanced Technologies. In this stage, breakthroughs in advanced technologies emerge, such as brain-computer interfaces [147, 81, 254], nanotechnology [136, 19, 61], and holographic technology [22, 273, 15]. These innovations enable unprecedented levels of interaction between humans and machines by augmenting human senses and cognition. For instance, intelligent chips [86, 115] may be developed that integrate directly with the human body, creating a seamless interface between biological and digital systems. Such advances also change the nature of the data. Unlike traditional static formats such as text, images, or videos, the data in this stage becomes more comprehensive, encompassing dynamic, real-time, and sensory-rich formats, such as neural activity patterns or real-time environmental data from augmented reality systems [31, 32]. In addition to the advancements in AI seen during the Data-Driven Era, this stage further expands the boundaries of human perception and cognition, revolutionizing how people experience and interact with the world. However, widespread application of these techniques is limited by energy constraints and the significant resources required for implementation.
Stage 4: Breakthroughs in Fundamental Theories. In this stage, humanity experiences a new wave of breakthroughs in fundamental theories, such as Einstein’s Relativity Theories [57, 58] and Quantum Mechanics [59, 279]. These theoretical advances redefine our understanding of the universe and offer new possibilities for innovations in technologies. Such breakthroughs may also unlock solutions to longstanding challenges in human history, i.e., the energy issues, offering clean, abundant, and sustainable energy sources.
@@ 4.3 Ultimate Level: Revolutionary Era
Stage 5: Energy Revolution. A new era arrives as humanity finally achieves energy revolution, unraveling the mysteries of sustainable power generation. This may be achieved through mature technologies that are currently on the horizon, such as breakthroughs in large-scale controlled nuclear fusion [70, 200, 183], or revolutionary solar technologies [23, 137], together with advancement in nanotechnologies [136, 19, 61] that produce new materials for collecting and storing solar energy efficiently. The energy issue may also be solved through innovations that are beyond the scope of today’s technologies, such as deploying satellites for gathering and storing solar energy in space. These satellites might then transmit the abundant energy to the receivers on Earth. Regardless of the specific technology, the flood of clean, abundant energy can eliminate the dependence on fossil fuels, and fundamentally revolutionize the global economic landscapes as well as human civilization.
Stage 6: The Internet of Everything. The final stage is the realization of Internet of Everything. Every object, living being, and system becomes part of a vast, intelligent network. This hyper-connected world is characterized by unprecedented levels of perception, communication, and interactions between humans and intelligent robots in an area. Language barriers may disappear, as human can perceive with chips for understanding, which enables ubiquitous real-time context-aware translations between people of different languages and cultures. Due to everyone is connected to a huge network, the boundary of human sense has been extended, which may also catalyze breakthroughs in holographic technology [15, 22, 273]. Remote experiences become true; working, entertainment, and communication are revolutionized, and human has entered the ultimate advanced era.
@@ 4.4 Evolving AI Safety Paradigms
Data-Driven Era. Current AI safety practices primarily focus on data safety and model safety [37, 102, 42, 220, 167, 193, 253, 225, 224]. A critical aspect involves data privacy and security, particularly given the vast amounts of sensitive information used in AI training [118, 120, 83, 117]. AI safety researchers and practitioners also explore adversarial attacks to training data or AI models [9, 35, 119, 244, 211, 269], as well as malicious inputs designed to manipulate AI systems or induce producing unsafe content, e.g., jailbreaking or prompt injection attacks [75, 292, 126, 33, 90, 258, 293]. Correspondingly, red-teaming ML models [12, 68, 173, 197, 243] is widely used to identify potential vulnerabilities in ML models with attacks to further enhance the robustness of AI models, and significant work has been proposed to enhance security and robustness of AI systems [27, 240, 290]. Additionally, researchers have been engaging in AI governance to enhance AI accountability [21], transparency and explainability [291] in AI systems to ensure the decision-making processes are understandable, accountable, and aligned with regulations. Other significant areas of focus is aligning AI outputs with human values [249, 174, 232, 206, 245].
@@
Neural Augmentation Era. Advancements in cognitive chips and brain-computer interfaces have extended the boundaries of human cognition and have changed the way of human perceiving the world. With the integration of intelligent chips and their direct neural connections to AI systems, users can interact with AI with high flexibility, and the customization of AI services towards different user groups or domain knowledge might be a critical focus [248]. Individuals might rely on AI in different ways based on their background and cognitive needs, thus, personalized experiences of AI services should be explored for different user needs and preferences.
@@
Revolutionary Era. In the revolutionary era, AI safety challenges may focus on the complex interactions between interconnected entities, such as humans and intelligent robots, within the network. As users may have diverse needs and robots will provide customized services, ensuring the safety and reliability of these interactions becomes challenging. Robots may need to request and integrate data from multiple external sources in real-time to fulfill user requests. Managing the quality and security of these interactions is crucial, as any breakdown or interference could lead to incorrect or even harmful outcomes. Ensuring seamless communication, data integrity, and decision-making in such a sophisticated network may require smarter, more efficient methods, such as advanced cache management systems, to support the dynamic and complex interactions. This requires addressing issues of latency, data accuracy, and the security of transmitted information to maintain a safe and reliable network to ensure high-quality interactions between humans and robots in the network.
@@
By aligning AI with the progress of human civilization, we can see the path of AI safety evolves in different stages, which reflects the growing complexity of the roles of AI in human society. Initially, the focus is on addressing safety issues related to data and models. As AI and novel technologies become more developed, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs. Finally, when entering the era of Internet of Everything, challenges in AI safety further evolve, focusing on complex interactions between humans and robots in the same network.
These insights offer guidelines for what we can do for AI safety at the current stage. In the following sections, we will overview today’s AI safety landscape, then reflect on the progress in this area so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed in the present stage.
@@ 5 Back to the 2020s: Today’s AI Safety Landscape
This section overviews mainstream AI safety practices so far, and project them against the blueprint future in §3.1 to guide our current efforts in AI safety.
@@ 5.1 Jailbreaking and Red Teaming
Jailbreaking [49, 123, 175, 207, 215, 219, 234, 250, 258, 268, 293] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs. Malicious users or adversaries utilize these techniques to manipulate the model into producing abnormal or harmful responses, such as giving instructions for illegal activities and generating offensive content. Common jailbreak methods include crafting inputs to trick the model, planting backdoors to exploiting vulnerabilities of the model, etc. Such activities pose significant risks by undermining safety measures of AI models, making it challenging to ensure the model to behaves appropriately across all scenarios.
@@
Red-teaming is a common practice for mitigating unsafe behaviors in LLMs, where teams of experts simulate attacks or misuse scenarios to identify potential vulnerabilities in AI models [68, 243, 197, 12, 173, 128, 275]. By actively challenging AI models and probing for weaknesses through adversarial inputs, bias exploitation, or system manipulation, it helps uncover potential risks that may not be evident in normal usage, such as biases, or failure modes that might lead to undesirable behaviors in AI models. The outputs from red-teaming can be used to train the model to be less likely to cause harm or steer it away from undesirable outputs. This proactive approach has proven effective in stress-testing AI models and has been widely used in industry practices [18, 173, 68].
@@
As AI systems grow more sophisticated, red teaming must continually adapt to keep pace with increasingly complex attacks, which calls for more advanced tools and interdisciplinary expertise [153]. Also, model owners can conduct regular and diverse red teaming to anticipate and mitigate risks in evolving use cases. Moreover, current red-teaming practices are often static, typically conducted after model pretraining and consisting of one-time attacks. While this approach provides useful insights, it does not fully account for the evolving user needs, especially when dealing with customized AI services. To address this, red teaming should simulate diverse user groups and scenarios, such as simulating a diverse range of user groups and their unique needs, to address the customization of AI services and enable a comprehensive evaluation of potential vulnerabilities across different contexts.
5.2 Content Moderation and Guardrails
@@
Content moderation involves detecting and reviewing user input and AI-generated content to ensure it complies with safety, ethical, and regulatory standards [133, 112, 80, 91]. Moderation approaches leverages rule-based methods, machine learning classifiers, and human oversight to monitor, evaluate, and manage inputs and outputs of AI models. These approaches identify and address inappropriate content, such as toxicity, bias, hallucinations, sensitive information, and jailbreaks, to ensure that the AI models do not produce or engage with harmful content, which maintains responsible AI behaviors.
Guardrails [141, 189] employ customizable AI workflows for ensuring the quality of the model outputs while enhancing the control of the whole workflows. It functions as an intermediary layer between users and AI models, enabling users to add components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
@@
As we reflect on practices for AI moderation and guardrails, several key challenges and considerations have emerged. The balance between utility and safety is crucial and may require dynamic solutions according to the changing contexts. Overly strict moderation can cause an AI system to be helpless, e.g., an LLM that always responds “sorry, I cannot answer this question” is completely safe but useless. Thus, defining appropriate thresholds for acceptable responses is challenging and requires careful consideration of the changing context and potential risks. Also, customization for moderation should be applied for different use cases and user groups, as certain information may be safe for one user group but inappropriate or dangerous for another. As an example, a chemistry industry practitioners may need specific and technical information on dangerous substances like explosives for professional purposes, while such information should be restricted for the general public.
@@
Guardrail - Guardrails AI [141], Nvidia Nemo Guardrails [189], Protect AI [220]
@@
Defense Methods [199, 88, 240, 176, 257, 262, 133, 29, 66, 247, 278, 88, 168, 74, 40, 51, 209, 208], etc OpenAI Rule-based rewards [143]; DeepMind Weight Averaged Reward Models [188]; others see Moderation, Guardrail, and Privacy
@@
AI governance [276, 277, 6, 104, 56], etc Giskard [42], ProtectAI [220], Calypso AI [26], Saidot [202], Arize [253], Dynamo AI [55], Credo AI [212], Google [187], DataGrail [193], Dastra [225], OneTrust [224], Relyance AI [99], Transcend [100], Zendata [229], DeepMind SynthID watermark [45, 47]

### v3

@@ 4.1 Rookie Level: Data-Driven Era
The data-driven era represents the earliest phase of AI development, where AI models rely heavily on the quality, availability, and quantity of training data [135, 261, 269]. AI makes probabilistic inferences based on patterns found in massive training datasets, thus failing to achieve genuine understanding of data [100, 245]. While AI in this stage cannot be regarded as real intelligence, it can still assist people to solve problems and enhances productivity in human society.
Stage 1: Past, Current, and the Near Future. This stage is characterized by developing diverse AI models to process various forms of data across different scenarios [135, 53, 66, 215, 236, 268]. Researchers have been engaging in refining model architectures and developing advanced algorithms, enabling more sophisticated analysis of data in different formats. The thrive of LLMs exemplifies this progress, demonstrating unprecedented capabilities of AI models in language processing and generation [268, 297, 236]. Concurrently, significant advancements have also been observed in other domains such as computer vision [105, 226, 264], image processing [241, 25, 221], video analysis [228, 270, 276], and audio recognition [67, 182, 4, 110]. This stage is driven by the exploration of vast datasets and the continuous refinement of algorithms, which consistently pushes the boundaries of AI capabilities across specific domains.
Stage 2: Interdisciplinary AI. This stage is characterized by the integration of AI through cross-disciplinary applications, driving significant innovations and transformations in traditional fields such as healthcare [236, 2], finance [268, 292], and education [188, 139]. In this stage, AI applications would unlock new opportunities across industries and reshape human society. Mature multi-agent systems [120, 253, 272, 82] that leverage AI entities with specialized capabilities might be widely deployed to assist people to solve sophisticated problems. Such agents will participate in decision-making processes or represent users to complete some tasks. For example, in healthcare, a multi-agent system may involve collaborations between multiple specialized medical agents and integrate data from various sources to formulate comprehensive treatment plans. The rise of interdisciplinary AI integration will enhance AI’s utility in increasingly complex scenarios. While AI at this stage does not yet achieve true intelligence, it becomes deeply integrated across various domains in human society and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
@@ 4.2 Champion Level: Neural Augmentation Era
In the neural augmentation era, humans may witness breakthroughs in fundamental theories and advanced technologies, which may extend the boundary of human cognition and enhance life experience.
Stage 3: Breakthroughs in Advanced Technologies. At this stage, breakthroughs in advanced technologies, such as brain-computer interfaces [84, 151, 262], nanotechnology [140, 19, 64], and holographic technology [22, 283, 15], augment human senses and cognition, enabling unprecedented levels of interactions between humans and machines. Language barriers may disappear, as human can perceive with chips for understanding, which enables ubiquitous real-time context-aware translations between people of different languages and cultures. Intelligent chips [89, 119] might be integrated into human bodies, creating a seamless interface between biological and digital systems. Such advancements may change the nature of the data. Unlike data of static formats such as text, images, or videos, the data might be more comprehensive, in dynamic and sensory-rich formats, such as neural activity patterns or real-time environmental data from augmented reality systems [31, 32]. This stage expands the boundaries of human perception and cognition, revolutionizing how people experience and interact with the world. However, applying these techniques will be still challenging due to the energy constraints and the significant resources required by these technologies.
Stage 4: Breakthroughs in Fundamental Theories. At this stage, humanity may witness a new wave of breakthroughs in fundamental theories, such as Einstein’s theories of relativity [60, 61] and quantum mechanics [62, 289]. These theoretical advancements have the potential to redefine our understanding of the universe, and, moreover, offer new possibilities for groundbreaking technological innovations and provide critical insights into solving the longstanding challenges in human history, i.e., the energy issues.
@@ 4.3 Ultimate Level: Revolutionary Era
Stage 5: Energy Revolution. A new era would arrive as humanity finally achieves energy revolution, unraveling the mysteries of sustainable power generation. This may be achieved through mature technologies that are currently on the horizon, such as breakthroughs in large-scale controlled nuclear fusion [73, 205, 187], or revolutionary solar technologies [23, 141], together with advancement in nanotechnologies [140, 19, 64] that produce new materials for collecting and storing solar energy efficiently. The energy issue may also be solved through innovations that are beyond the scope of today’s technologies, such as deploying satellites for gathering and storing solar energy in space. These satellites might then transmit the abundant energy to the receivers on Earth. Regardless of the specific technology, the flood of clean, abundant energy can eliminate our dependence on fossil fuels, and revolutionize the global economic landscapes as well as human civilization.
Stage 6: The Internet of Everything. The final stage is the Internet of Everything. Every object, living being, and system become parts of a vast, intelligent network. This hyper-connected world is characterized by unprecedented levels of perception, communication, and interactions between humans and intelligent robots. In this era, everyone is connected to a huge network, and the boundary of human sense has been extended, which may also catalyze breakthroughs in holographic technology [15, 22, 283]. Remote experiences become true; working, entertainment, and communication are revolutionized, and human has entered the ultimate advanced era.
@@ 4.4 Evolving AI Safety Paradigms
Data-Driven Era. Current AI safety practices primarily focus on data safety and model safety [39, 106, 44, 225, 171, 198, 260, 230, 229]. AI safety researchers and practitioners have explored adversarial attacks targeting training data and AI models [9, 36, 123, 250, 216, 279] to achieve malicious goals, for example, designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection attacks [78, 303, 130, 34, 93, 266, 304]. Correspondingly, red-teaming ML models [12, 71, 177, 202, 249] is widely used to identify potential vulnerabilities in AI models, and significant work has been proposed to enhance security and robustness of AI systems [27, 246, 301], aligning AI with human values [256, 178, 237, 211, 252]. Data privacy is also a critical aspect, given the vast amounts of sensitive information in the training data [122, 124, 86, 121]. Additionally, researchers have been engaging in AI governance to enhance AI accountability [21], transparency and explainability [302] in AI systems to ensure the behaviors of AI are understandable, accountable, and aligned with regulations [21, 136, 219]. .
@@
Neural Augmentation Era. Cross-disciplinary AI applications might have been widely deployed in human society, and advancements in intelligent chips [89, 119] and brain-computer interfaces [84, 151, 262] might extend the boundaries of human cognition and have changed the way of human perceiving the world. With the integration of intelligent chips and their neural connections to intelligent systems, users can interact with machines and AI with high flexibility, and the customization of AI services for different user needs might be a critical focus [255]. Individuals might rely on AI in different ways, and personalized experiences of AI services might be explored based on users’ background and requirements.
@@
Revolutionary Era. In the revolutionary era, AI safety challenges may focus on complex interactions between interconnected entities, e.g., humans and intelligent robots, within the network. Robots might provide customized services according to different user requirements, and ensuring the safety and reliability of the interactions between them becomes challenging. Robots may need to integrate data from external sources in real-time to fulfill user requests. Managing the quality and security of these interactions is crucial, as any breakdown or interference could lead to incorrect or even harmful outcomes. Thus, ensuring seamless communication, data integrity, and efficient decision-making in such a sophisticated network may require smarter and more efficient methods for managing, such as advanced cache management systems, to support the dynamic and complex interactions between humans and robots in the network.
@@
By aligning advancements of AI with human civilization, we can see the focuses of AI safety evolve in different stages. Initially, the focus is on addressing safety issues related to data and models. With progress of AI and advanced technologies, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs. Finally, when humanity entering the era of Internet of Everything, challenges in AI safety further evolve, focusing on complex interactions between humans and robots in the network. These insights offer guidelines for our efforts on AI safety at the current stage. The following sections will overview today’s AI safety landscape, then reflect on the progress in this area so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed at the current stage.
@@ 5 Back to the 2020s: Today’s AI Safety Landscape
This section overviews current mainstream AI safety practices, and project them against the blueprint future in §3.1 to guide our current efforts on AI safety.
@@ 5.1 Jailbreaking and Red Teaming
Jailbreaking [51, 127, 179, 212, 220, 224, 240, 257, 266, 277, 304] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs. Malicious users or adversaries can exploit vulnerabilities in AI models through crafted inputs or backdoors, inducing them to produce inappropriate responses, such as giving instructions for illegal activities or generating offensive content. These activities pose significant risks by compromising safety measures of AI models, making it challenging to ensure the model to behave appropriately, especially in real-world scenarios where the potential inputs are unpredictable.
@@
Red-teaming is a common practice that simulates attacks or misuse scenarios to identify vulnerabilities in AI models, thereby mitigating unsafe behaviors [71, 249, 202, 12, 177, 132, 285]. It challenges AI models with adversarial inputs, bias exploitation methods, and system manipulation to uncover potential risks, e.g., hidden biases or failure modes, that may not be evident during normal usage. The outcomes from red-teaming can be leveraged to refine AI models, reducing undesired behaviors and inappropriate outputs. This approach is effective in stress-testing AI models and has been widely employed in industry [18, 177, 71].
@@
As AI systems grow more sophisticated, red-teaming must continually adapt to counter increasingly complex attacks, which calls for more advanced tools and interdisciplinary expertise [157]. Model owners can conduct diverse red teaming regularly to anticipate and mitigate risks in evolving use cases. Also, current red-teaming practices are often static, typically conducted after model pretraining with one-time attacks. Thus, these practices potentially fail to address evolving user needs and complex interactions between users and models, especially when dealing with customized AI services for different user groups. To address this, red teaming should involve more diverse use cases, such as employing different user groups and simulating complex interactions with dynamic contexts, to address safety risks in customized AI services and enable a comprehensive evaluation of potential vulnerabilities across different scenarios.
5.2 Content Moderation and Customized Moderation Pipeline
@@
Content moderation involves detecting and reviewing both user input and AI-generated content to ensure compliance with safety, ethical, and regulatory standards [137, 116, 83, 94]. Moderation approaches employ rule-based methods, machine learning classifiers, and human oversight to monitor, evaluate, and manage the inputs and outputs of AI models. These approaches identify inappropriate content, such as toxicity, bias, hallucinations, private information, and jailbreaks [251]. By preventing AI models from engaging with or producing harmful content, content moderation plays a critical role in maintaining responsible AI behavior. The customized moderation pipeline [145, 194] employs multiple components at different stages of LLM inference to enhance the quality of the model outputs while enhancing the control of the whole AI workflows. It functions as a flexible intermediate layer between users and AI models, enabling users to add customized components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
@@
As we reflect on practices for AI moderation and customized moderation pipelines, several key challenges and considerations have emerged. 1) Hallucination remains a significant issue and is unlikely to solve completely, regardless of how AI evolves. That is because hallucinations stem from nearly every stage of the LLM lifecycle, including training data quality, model architecture, and inference processes, etc [92]. Despite advancements in AI technology, hallucinations persist as an inherent challenge that requires constant attention in moderation practices. 2) The balance between utility and safety is crucial and may require different solutions according to the changing contexts and user needs. Overly strict moderation can cause an AI system to be helpless, e.g., an LLM that always responds “sorry, I cannot answer this question” is completely safe but useless. Defining appropriate thresholds for acceptable responses is challenging and requires considering potential risks and the context of user queries. 3) Customization of moderation methods should be applied for different use cases, as certain information may be safe for one user group but inappropriate for another. As an example, a chemistry industry practitioners may need specific and technical information on dangerous substances like explosives for professional purposes, while such information should be restricted for the general public. Current moderation practices might fail to include such cases, and applying identical restrictions between different user expertise or context might limit the utility of AI in specialized fields. 4) Some moderation measures, such as detection of toxicity, bias, stereotypes, out-of-distribution content, might be unnecessary for most real-world use cases. This is because most of the current AI applications and multi-agent systems [120, 37] leverage existing LLM APIs [139, 188, 137], open-source LLMs [239, 278, 97], or utilizing opensource LLMs as base models, many of which already have built-in components or have done safety alignments for managing inappropriate inputs and outputs. Incorporating certain detection mechanisms may offer limited benefits to the safety of LLM inputs and outputs while increasing unnecessary computational overhead.
@@
Customized Workflow - Guardrails AI [145], Nvidia Nemo Guardrails [194], Protect AI [225]
@@
Defenses [204, 91, 246, 180, 265, 271, 137, 29, 69, 254, 288, 91, 172, 77, 42, 54, 214, 213], etc OpenAI Rule-based rewards [147]; DeepMind Weight Averaged Reward Models [192]; others see Moderation, Guardrail, and Privacy
@@
AI governance [286, 287, 6, 108, 59, 33], etc Giskard [44], ProtectAI [225], Calypso AI [26], Saidot [207], Arize [260], Dynamo AI [58], Credo AI [217], Google [191], DataGrail [198], Dastra [230], OneTrust [229], Relyance AI [103], Transcend [104], Zendata [234], DeepMind SynthID watermark [47, 49]